# WoCoCo: Learning Whole-Body Humanoid Control with Sequential Contacts

### v1

@@ Abstract
Humanoid activities involving sequential contacts are crucial for complex robotic interactions and operations in the real world and are traditionally solved by model-based motion planning, which is time-consuming and often relies on simplified dynamics models. Although model-free reinforcement learning (RL) has become a powerful tool for versatile and robust whole-body humanoid control, it still requires tedious task-specific tuning and state machine design and suffers from long-horizon exploration issues in tasks involving contact sequences. In this work, we propose WoCoCo (Whole-Body Control with Sequential Contacts), a unified framework to learn whole-body humanoid control with sequential contacts by naturally decomposing the tasks into separate contact stages. Such decomposition facilitates simple and general policy learning pipelines through task-agnostic reward and sim-to-real designs, requiring only one or two task-related terms to be specified for each task. We demonstrated that end-to-end RL-based controllers trained with WoCoCo enable four challenging whole-body humanoid tasks involving diverse contact sequences in the real world without any motion priors: 1) versatile parkour jumping, 2) box loco-manipulation, 3) dynamic clap-and-tap dancing, and 4) cliffside climbing. We further show that WoCoCo is a general framework beyond humanoid by applying it in 22-DoF dinosaur robot loco-manipulation tasks.
@@ 1 Introduction
In comparison, a single model-based solver such as Crocoddyl [5] can address multiple tasks with different contact sequences, requiring only slight adjustments to costs (few intuitive task-related terms). The question is: How can we achieve such simplicity and adaptability with an RL framework? To this end, we identify three primary challenges, which together lead to hard exploration: (1) Contacts are sparse, especially when coupled with other whole-body motion goals such as balancing and posture maintenance; (2) Robots may avoid exploring the whole long horizon due to the compounding risks; and (3) Sim-to-real transfer is non-trivial, and is often achieved through domain randomization [27] and task-irrelevant regularization rewards, which can hinder exploration.
In this work, we address these challenges by proposing WoCoCo, a general RL framework for whole-body humanoid control with sequential contacts. In WoCoCo, we reformulate the problem as the sequential fulfillment of multiple contact stages (detailed in Section 2), which breaks down the exploration burden into separate stages. This transforms the challenges to three key questions: Q1: How to reach desired contact states within each stage? Q2: How to streamline exploration across multiple contact stages? and Q3: How to develope a compatible sim-to-real pipeline?
@@ 2 Overview: Learning with Sequential Contacts and Task Decomposition
3 How to Train Your Droid
@@ 3.1 WoCoCo Rewards
In this work, we aim to define curiosity rewards with task-agnostic observations that can be redundant and high-dimensional. We find that RND can over-explore states that do not generate meaningful behaviors, similar to the ”DANCING WITH SKULLS” behaviors reported in the RND work [36]. Instead, we propose to use count-based curiosity rewards via random neural network (NN) based hash, inspired by Tang et al. [37] and Charikar [38].
@@ 4.2 Case II: Anywhere-to-Anywhere Box Loco-Manipulation
Box loco-manipulation is an important application of humanoids and has been well studied with model-based controllers [45]. However, model mismatch and perturbations such as uneven terrains pose significant challenges to these controllers, for which RL has proven to be a promising solution [13, 22].
@@ 6 Limitation and Future Works
One limitation of our work is the lacking knowledge of when the controller will fail. In contrast, model-based methods can explicitly tell whether they can find a feasible solution. Therefore, we may explore failure predictors [56] and other safety assessment methods in the future [57].
@@ Appendix B Ablation Baseline Behaviors
We visualize the learned behaviors of baselines in ablation studies in Figure 8.
@@ Appendix H Policy Observations and Architechtures
The policy observations in this work consist of two parts for all policies: proprioception and goal representations. Exteroception is not used here but our framework does not exclude it.

### v2

@@ Abstract
Humanoid activities involving sequential contacts are crucial for complex robotic interactions and operations in the real world and are traditionally solved by model-based motion planning, which is time-consuming and often relies on simplified dynamics models. Although model-free reinforcement learning (RL) has become a powerful tool for versatile and robust whole-body humanoid control, it still requires tedious task-specific tuning and state machine design and suffers from long-horizon exploration issues in tasks involving contact sequences. In this work, we propose WoCoCo (Whole-Body Control with Sequential Contacts), a unified framework to learn whole-body humanoid control with sequential contacts by naturally decomposing the tasks into separate contact stages. Such decomposition facilitates simple and general policy learning pipelines through task-agnostic reward and sim-to-real designs, requiring only one or two task-related terms to be specified for each task. We demonstrated that end-to-end RL-based controllers trained with WoCoCo enable four challenging whole-body humanoid tasks involving diverse contact sequences in the real world without any motion priors: 1) versatile parkour jumping, 2) box loco-manipulation, 3) dynamic clap-and-tap dancing, and 4) cliffside climbing. We further show that WoCoCo is a general framework beyond humanoid by applying it in 22-DoF dinosaur robot loco-manipulation tasks. Website: lecar-lab.github.io/wococo/.
@@ 1 Introduction
In comparison, a single model-based solver such as Crocoddyl [5] can address multiple tasks with different contact sequences, requiring only slight adjustments to costs (few intuitive task-related terms). The question is: How can we achieve such simplicity and adaptability with an RL framework? Besides, regarding effective policy learning, we also identify three challenges: (1) Contacts are sparse, especially when coupled with other whole-body motion goals such as balancing and posture maintenance; (2) Robots may avoid exploring the whole long horizon due to the compounding risks; and (3) Sim-to-real transfer is non-trivial, and is often achieved through domain randomization [27] and task-irrelevant regularization rewards, which can hinder exploration.
In this work, we address the question by proposing WoCoCo, a general RL framework for whole-body humanoid control with sequential contacts. In WoCoCo, we reformulate the problem as the sequential fulfillment of multiple contact stages (detailed in Section 2), which also breaks down the exploration burden into separate stages. This then transforms each challenge to a question: Q1: How to reach desired contact states within each stage? Q2: How to streamline exploration across multiple contact stages? and Q3: How to develope a compatible sim-to-real pipeline?
@@ 2 Overview: Learning with Sequential Contacts and Task Decomposition
3 WoCoCo Rewards and Sim-to-Real Transfer
@@ 3.1 WoCoCo Rewards
In this work, we aim to define curiosity rewards with task-agnostic observations that can be redundant and high-dimensional. We find that RND can over-explore states that do not generate meaningful behaviors, similar to what is reported in the RND work’s Section 3.7 [36]. Instead, we propose to use count-based curiosity rewards via random neural network (NN) based hash, inspired by Tang et al. [37] and Charikar [38].
@@ 4.2 Case II: Anywhere-to-Anywhere Box Loco-Manipulation
Box loco-manipulation is an important application of humanoids and has been well studied with model-based controllers [45]. However, model mismatch and perturbations such as uneven terrains pose significant challenges to these controllers, for which RL can be a promising solution [13, 22].
@@ 6 Limitation and Future Works
One limitation of our work is the lacking knowledge of when the controller will fail. In contrast, model-based methods can explicitly tell whether they can find a feasible solution. Therefore, we may explore failure predictors [56] and other safety assessment methods in the future [57]. Besides, if the contact sequence length is unknown a priori, we may need heuristic reward clamping to avoid the robot exploiting the stage count reward.
@@ Appendix B Ablation Baseline Behaviors
We visualize the learned behaviors of baselines in ablation studies in Figure 8. We also show their corresponding learning curves below for the average curiosity values and task progress (i.e., the number of fulfilled contact stages divided by its maximum number).
@@
Supplementary Figure 9: Corresponding learning curves of baselines. The convergent average progress of WoCoCo is also visualized in red, marking successful policy learning.
@@ Appendix H Policy Observations and Architechtures
The policy observations in this work consist of two parts for all policies: proprioception and goal representations. Camera and LiDAR observations are not used here but our framework does not exclude them. The goal representations are provided by the motion capture system as a prototype.
@@ H.2 Goal Representations
For versatile parkour jumping, we feed the future two stages’ goal representations to the policy network, so the robot can adapt the foothold for future goals. For other tasks, we feed only the current stage’s goal representations. This design allows our policies to generalize to varying sequence lengths during deployment.
@@ I.2 NN Architechtures
Appendix K Significance of Sim-to-Real Curriculum
Our training curriculum for sim-to-real transfer has 3 phases. Here we show the significance of such design by ablations.
K.1 Moving Domain Randomization to Phase 1
If we move the domain randomization from phase 2 to phase 1, i.e., to merge the first two phases, we find the policy learning can be hindered by heavy randomization. For example, when learning the parkour jumping policy, this merge leads to the average progress converging below $0.6$, compared to the ones $>0.75$ in standard WoCoCo.
K.2 Canceling Phase 3
In phase 3, we graudally increase the weights of the regularization terms. Without this phase, we find the learned behaviors can be aggressive in the real world, which may lead to failures even for the least dynamic task, i.e., cliffside climbing, as shown in Figure 11.
Supplementary Figure 11: Without the last phase of the sim-to-real curriculum, the real-world behaviors can be jerky. For example, during cliffside climbing, the robot often stepped out of the cliff edge (indicated by white dotted lines) due to jerky motions, while all of the contact goals were set within the boundary.
Appendix L Deployment Details
L.1 Control
Our policy updates at 50 Hz, and the PD controller updates torque commands at 200 Hz. We apply a Butterworth low-pass filter of 4-Hz bandwidth to the PD targets to avoid jerky outputs.
L.2 Real-to-Sim
We found the official URDF file contains significantly biased torso mass and wrong foot geometry, so we weighed the torso mass and measured the foot geometry by our own.